Pioneer Chicken advertises​ "lite" chicken with​ 30% fewer calories than standard chicken. When the process for​ "lite" chicken breast production is in​ control, the average chicken breast contains 450 calories, and the standard deviation in caloric content of the chicken breast population is 20 calories.Pioneer wants to design an x​-chart to monitor the caloric content of chicken​ breasts, where 25 chicken breasts would be chosen at random to form each sample.

​a) What are the lower and upper control limits for this chart if these limits are chosen to be four standard deviations from the​ target?

Upper Control Limit ​(UCLx​)= ? calories ​(enter your response as an​ integer).

Lower Control Limit ​(LCLx​) = ? calories ​(enter your response as an​ integer).

​b) What are the limits with three standard deviations from the​ target?

Upper Control Limit ​(UCLx​) = ? calories ​(enter your response as an​ integer).

Lower Control Limit ​(LCLx​) ​= ? calories ​(enter your response as an​ integer).
Answer #1

1. FORMULA
CONTROL LIMIT = X +- (Z * STDEV / SQRT(N)

Where X = mean
Z = sigma
STDEV = standard deviation
N = number of samples

VALUES:

X = 450
Z = 4
STDEV = 20
N = 25

UCL = 450 + (4 * 20 / SQRT(25)) = 466
LCL = 450 - (4 * 20 / SQRT(25)) = 434

2. FOR Z = 3

FORMULA
CONTROL LIMIT = X +- (Z * STDEV / SQRT(N)

Where X = mean
Z = sigma
STDEV = standard deviation
N = number of samples

VALUES:

X = 450
Z = 3
STDEV = 20
N = 25

UCL = 450 + (3 * 20 / SQRT(25)) = 462
LCL = 450 - (3 * 20 / SQRT(25)) = 438